OpenStudy (anonymous):

tom the cat is chasing jerry the mouse across a table surface 1.5m off the floor. Jerry steps out of the way at the last second and tom slides off the edge of the table at a speed of 5.7m/s. where will tom strike the floor?

OpenStudy (anonymous):

Hola. What variable(s) do you need to solve this?

OpenStudy (anonymous):

Distance? Velocity? Time? Density of a cat? Purrrmitivity of freespace?

OpenStudy (anonymous):

Time. You have velocity in the x direction, acceleration, and the height of the table. You can find the time - how long Tom is in the air for - by seeing how long it takes him to hit the ground; aka, how long it takes him to travel 1.5m from rest while being accelerated 9.8 m/s^2 (gravity). All you need is the first equation of motion; \[ h = \cancel{v_{oy} t}^0 + \frac{1}{2} a t^2 \\ t = \sqrt{\frac{2h}{a}} \] Then the velocity Tom travels in the x direction - how far away from the table he hits the ground, is \[ x = vt \]